OpenStudy (anonymous):

A cube is painted red on all six sides, then cut into 125 smaller congruent cubes. A small cube is selected at random and rolled. What is the probability that it has a red side facing up?

OpenStudy (anonymous):

I guess you didn't understand what I said? Want to try again?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What were the numbers we had?

OpenStudy (anonymous):

none;1;2,3 sides are 27 cubes, 54 cubes, 36 cubes, and 8 cubes

OpenStudy (anonymous):

OK. So the Prob that you have a cube that's red on 3 sides is 8/125 and the Prob that when you throw it, it lands with red up is 3/8 so that's a total of 3/125.

OpenStudy (anonymous):

Do you see why that is?

OpenStudy (anonymous):

3/6

OpenStudy (anonymous):

8/125 3/6 = 4/125

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

This is what's called enumerating the sample space. You think of the possibilities, figure out the P of each, and add them up.

OpenStudy (anonymous):

Total P (of getting a red side up) = 8/125 x 3/6 + 36/125 x 2/6 + 54/125 x 1/6

OpenStudy (anonymous):

which equals 1/5

OpenStudy (anonymous):

Let's see 4 + 12 + 54 / 125 = 70 / 125?

OpenStudy (anonymous):

OK?

OpenStudy (anonymous):

yes